# Flexible Tool Selection through Low-dimensional Attribute Alignment of Vision and Language

### v2

@@ V Discussion and Conclusion
Our work establishes a cognitively inspired computational framework for flexible tool selection that achieves 74% accuracy while using significantly fewer parameters than state-of-the-art multimodal models. By formalizing the role of low-dimensional attribute space as a bridge between visual tool perception and linguistic task understanding, we demonstrate both the computational efficiency and cognitive plausibility of attribute-based tool selection.
The performance gap between visual (96.05%) and language (74.34%) pathways reveals a fundamental asymmetry in how attributes are extracted versus inferred: visual systems can directly extract explicit physical properties from images, while language systems must infer implied requirements from natural language descriptions. This asymmetry aligns with cognitive research showing that physical object properties are more directly accessible than functional requirements inferred from task contexts[70, 71]. Our ablation studies provide crucial insights into the functional primitives underlying tool selection, revealing that manipulation-related attributes (graspability, hand-relatedness, elongation) consistently prove most critical across modalities. This consistency supports the technical reasoning hypothesis in cognitive science, which emphasizes the importance of functional and physical property reasoning in tool use[72, 73].
@@
While our framework demonstrates significant advances, several limitations warrant consideration and suggest directions for future work. First, despite being theoretically motivated and empirically validated, our 13-dimensional attribute space represents a simplification of human tool cognition. These dimensions capture essential physical and functional properties but may not encompass all factors relevant to real-world tool selection, such as availability, cost, or contextual appropriateness[76]. Second, the visual-language performance gap suggests room for improvement in extracting attribute requirements from natural language descriptions. Third, our evaluation focuses on static tool selection rather than dynamic manipulation[77], which would require additional considerations of temporal sequences and motor control. Future research should explore expanding the attribute space to incorporate additional dimensions (e.g., material properties, temporal constraints) while investigating hierarchical or learned attribute spaces through unsupervised methods to discover optimal representations for specific domains. Extending the framework to dynamic tool manipulation and sequential task planning would better capture the full complexity of human tool use and its neural mechanisms.
This work advances our understanding of flexible tool selection through a computationally efficient and cognitively plausible framework that bridges visual perception and linguistic understanding. By demonstrating that attribute-based representations enable effective cross-modal matching for tool selection, we contribute to both cognitive science and computational modeling of human-like intelligent systems. Our approach provides a foundation for developing more interpretable, efficient, and neurally-grounded systems that reflect the remarkable flexibility of human tool use.

### v4

@@ V Discussion and Conclusion
Our work establishes a cognitively inspired computational framework for flexible tool selection that achieves 74% accuracy on standard benchmarks while using significantly fewer parameters than state-of-the-art multimodal models. Through evaluation across multiple dimensions—including model-human alignment analysis and novel tool generalization —we demonstrate both the computational efficiency and cognitive plausibility of attribute-based tool selection.
The performance gap between visual (96.05%) and language (74.34%) pathways reveals a fundamental asymmetry in how attributes are extracted versus inferred: visual systems can directly extract explicit physical properties from images, while language systems must infer implied requirements from natural language descriptions. This asymmetry aligns with cognitive research showing that physical object properties are more directly accessible than functional requirements inferred from task contexts[70, 71]. Our ablation studies provide crucial insights into the functional primitives underlying tool selection, revealing that manipulation-related attributes (graspability, hand-relatedness, elongation) consistently prove most critical across modalities. This consistency supports the technical reasoning hypothesis in cognitive science, which emphasizes the importance of functional and physical property reasoning in tool use[72, 73]. The minimal impact of psychological attributes (valence, arousal) suggests that affective properties, while potentially relevant in real-world contexts, are less critical for core tool selection decisions than functional characteristics.
@@
Our novel tool generalization experiment reveals both the promise and limitations of attribute-based approaches for handling unseen tool categories. The 22.6% Top-1 accuracy on completely novel tools, while exceeding chance performance (10%), indicates that learned attribute relationships enable some transfer to new categories. However, the substantial variation across individual novel tools (ranging from well above chance to below chance performance) suggests that generalization success depends heavily on the similarity between novel tool attributes and those represented in the training dataset.
While our framework demonstrates significant advances, several limitations warrant consideration and suggest directions for future work. First, despite being theoretically motivated and empirically validated, our 13-dimensional attribute space represents a simplification of human tool cognition. These dimensions capture essential physical and functional properties but may not encompass all factors relevant to real-world tool selection, such as availability, cost, or contextual appropriateness[76]. Second, the visual-language performance gap suggests room for improvement in extracting attribute requirements from natural language descriptions. Third, our evaluation focuses on static tool selection rather than dynamic manipulation[77], which would require additional considerations of temporal sequences and motor control. Fourth, the moderate model-human correlation (r = 0.373) suggests our attribute space may miss important factors in human decision-making, such as tool availability, safety concerns, or personal familiarity. Fifth, our generalization experiment reveals a fundamental limitation: while maintaining competitive performance with large multimodal models on known tools (74% vs 72-73%), our approach shows substantial performance drops on novel tool categories (22.6% vs 91.2% Top-1 accuracy for Gemini-2.0-Pro). This disparity stems primarily from training data limitations rather than architectural constraints—our visual encoder trained on 115 tool categories cannot match the recognition capabilities of models trained on millions of diverse tool images. Future research should explore expanding the attribute space to incorporate additional dimensions (e.g., material properties, temporal constraints) while investigating hierarchical or learned attribute spaces through unsupervised methods to discover optimal representations for specific domains. Extending the framework to dynamic tool manipulation and sequential task planning would better capture the full complexity of human tool use and its neural mechanisms.
This work advances our understanding of flexible tool selection through a computationally efficient and cognitively plausible framework that bridges visual perception and linguistic understanding. By demonstrating that attribute-based representations enable effective cross-modal matching for tool selection—while revealing important limitations in human alignment and novel tool generalization—we contribute to both cognitive science and computational modeling of human-like intelligent systems. Our evaluation across standard performance, human comparison, and generalization scenarios provides a foundation for developing more interpretable, efficient, and neurally-grounded systems that reflect both the capabilities and limitations of human tool use. Future advances in attribute space design and generalization methods will be crucial for deploying such systems in real-world applications where tool diversity and contextual complexity exceed current capabilities.
Code Availability
The complete source code for our attribute-based tool selection framework and the datasets are available at: https://drive.google.com/drive/folders/179W5KxPMzidjgg4hTzObrz-RvnXcVW0q?usp=sharing. Pre-trained model weights for all configurations tested in this paper are available at: https://1024terabox.com/s/17IdOOZ12GVtwJiWKySkOeQ